### v1

# Jarník-type theorem for self-similar sets

Abstract.
Let $K\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular measure $\mu$. For any $\tau>1/d$ and any ‘inhomogeneous’ vector ${\bm{\theta}}\in\mathbb{R}^{d}$, let $W_{d}(\psi_{\tau},{\bm{\theta}})$ denote the set of $(\psi_{\tau},{\bm{\theta}})$-well approximable numbers, where $\psi_{\tau}(q)=q^{-\tau}$. Assuming a local estimate for the $\mu$-measure of the intersections of $K$ with the neighborhoods of ‘rational’ vectors $(\mathbf{p}+{\bm{\theta}})/q$, we establish a sharp upper bound for the Hausdorff dimension of $K\cap W_{d}(\psi_{\tau},{\bm{\theta}})$, together with some nontrivial lower bounds when $\tau$ is below a certain threshold. One of the lower bounds becomes sharp in the one-dimensional homogeneous case ($d=1,\theta=0$) for a class of sufficiently thick self-similar sets $K$, and moreover $K\cap W_{1}(\psi_{\tau},0)$ has full $(\delta+\frac{2}{1+\tau}-1)$-Hausdorff measure. These results have several applications:
  1. (1)
    the set of homogeneous very well approximable numbers has full Hausdorff dimension within strongly irreducible self-similar sets in $\mathbb{R}^{d}$, extending a recent result of Chen [arXiv: 2510.17096];
  2. (2)
    the set of inhomogeneous very well approximable numbers has full Hausdorff dimension within sufficiently thick missing digits sets in $\mathbb{R}$, affirmatively answering a question posed by Yu [arXiv: 2101.05910].
Our applications build on the seminal works of Yu [arXiv: 2101.05910] and Bénard, He and Zhang [arXiv:2508.09076].
Key words and phrases:
Metric Diophantine approximation, self-similar sets, inhomogeneous approximation, Hausdorff measure
2020 Mathematics Subject Classification:
11J83, 28A80, 11K55, 28A78.
1. Introduction
Let $d\geq 1$ and $\mathcal{L}_{d}$ be the $d$-dimensional Lebesgue measure. We write $|\cdot|$ for the standard supremum norm in $\mathbb{R}^{d}$. For any $\mathbf{x}\in\mathbb{R}^{d}$, let $\|\mathbf{x}\|$ denote the distance from $\mathbf{x}$ to the nearest integer vector in $\mathbb{Z}^{d}$. Let $\psi:\mathbb{N}\rightarrow\mathbb{R}_{+}$ be a monotonic function and ${\bm{\theta}}\in\mathbb{R}^{d}$. The set of $(\psi,{\bm{\theta}})$-well approximable numbers is defined by
$W_{d}(\psi,{\bm{\theta}}):=\{\mathbf{x}\in\mathbb{R}^{d}:\|q\mathbf{x}-{\bm{\theta}}\|<\psi(q)\text{ for infinitely many }q\in\mathbb{Z}\}.$
If $\psi(q)=q^{-\tau}$, then we write $\psi_{\tau}$ in place of $\psi$. Since $W_{d}(\psi,{\bm{\theta}})$ is invariant under integer translations, we may restrict our attention to $\mathbf{x}\in[0,1]^{d}$. The metric properties of the set $0,1]^{d}\cap W_{d}(\psi,{\bm{\theta}})$ have been extensively studied in Diophantine approximation. A fundamental result in this direction is Khintchine’s theorem, which provides a zero–one law for the Lebesgue measure of $0,1]\cap W_{1}(\psi,0)$ in terms of the convergence or divergence of a natural volume sum.
Khintchine’s theorem ([21]). Let $\psi:\mathbb{N}\to\mathbb{R}_{+}$ be a monotonic function. Then,
$\mathcal{L}_{1}\big([0,1]\cap W_{1}(\psi,0)\big)=\begin{cases}0&\sum_{q=1}^{\infty}\psi(q)<\infty,\\[4.30554pt] 1&\sum_{q=1}^{\infty}\psi(q)=\infty.\end{cases}$
While Khintchine theorem gives a complete description of the Lebesgue measure of $0,1]\cap W_{1}(\psi,0)$, it does not capture the finer geometric structure of such set in the null case. In order to quantify the size of $0,1]\cap W_{1}(\psi,0)$ beyond Lebesgue measure, it is natural to consider its Hausdorff dimension or measure.
A seminal result in this direction is Besicovitch–Jarník’s theorem, which determines the Hausdorff dimension of the sets of well approximable numbers for monotonic functions.
Besicovitch–Jarník’s theorem ([5] and [18]). Let $\psi:\mathbb{N}\to\mathbb{R}_{+}$ be a monotonic function. Then,
$\operatorname{\dim_{H}}\big([0,1]\cap W_{1}(\psi,0)\big)=\frac{2}{1+\tau(\psi)},$
where
$\tau(\psi):=\liminf_{q\to\infty}\frac{-\log\psi(q)}{q}$
is the lower order of $\psi$ at infinity.
The above dimension result was further extended by Jarník [19] to a statement about the Hausdorff measure of well approximable sets. More precisely, one can determine the $s$-dimensional Hausdorff measure of $0,1]\cap W_{1}(\psi_{\tau},0)$ for all $s>0$. In this paper, we will adopt a special case of Jarník’s theorem that suffices for our purposes.
Jarník’s theorem ([19]). Let $\tau>1$. Then,
$\mathcal{H}^{s}\big([0,1]\cap W_{1}(\psi_{\tau},0)\big)=\begin{cases}0&\displaystyle s>\frac{2}{1+\tau},\\[4.30554pt] \infty&\displaystyle s\leq\frac{2}{1+\tau}.\end{cases}$
Jarník’s theorem turns out to be very useful in describing the Hausdorff measure of the sets of numbers with a given irrationality exponent.
For a point $\mathbf{x}\in\mathbb{R}^{d}$, its irrationality exponent $\tau(\mathbf{x})$ is defined by
$\tau(\mathbf{x}):=\sup\big\{\tau>0:\mathbf{x}\in W_{d}(\psi_{\tau},0)\big\}.$
With this definition in hand, we define the set of numbers with irrationality exponent equal to a given value $\tau$ by
$E_{d}(\tau,\mathbf{0}):=\big\{\mathbf{x}\in\mathbb{R}^{d}:\tau(\mathbf{x})=\tau\big\}.$
A corollary of Jarník’s theorem. Let $\tau>1$. Then,
$\mathcal{H}^{s}\big([0,1]\cap E_{1}(\tau,0)\big)=\begin{cases}0&\displaystyle s>\frac{2}{1+\tau},\\[4.30554pt] \infty&\displaystyle s\leq\frac{2}{1+\tau}.\end{cases}$
For a proof of this corollary, see [4]. These one-dimensional homogeneous results, which provide a foundation for understanding the metric properties of well-approximable sets, have also been extended to higher dimensions $d\geq 2$ and to the inhomogeneous setting $\bm{\theta}\neq\bm{0}$ [29, 1, 8, 7, 4].
In recent years, increasing attention has been paid to the study of intersections of $W_{d}(\psi,{\bm{\theta}})$ with fractal sets. This line of investigation was motivated by a question of Mahler [27]:
Question (Mahler, [27]). How closely can irrational elements of Cantor’s set be approximated by rational numbers?
Although Mahler himself did not pose a more specific question, this nonetheless inspired subsequent studies on the metric properties of the intersection of $W_{d}(\psi,{\bm{\theta}})$ with self-similar sets, of which the Cantor set is a special case. Recent years have witnessed significant progress in this direction, particularly with the establishment of Khintchine-type theorems for self-similar measure and Besicovitch–Jarník-type theorem for missing digits sets. In the following, we discuss the problem in two parts, focusing separately on the Lebesgue measure and the Hausdorff dimension.
For the Lebesgue measure, Kleinbock, Lindenstrauss and Weiss [22] established that if $\mu$ is a friendly measure—which includes strongly irreducible self-similar measures—then $\mu(W_{d}(\psi_{\tau},\mathbf{0}))=0$ for all $\tau>1/d$. These results were later extended by Pollington and Velani [28] to arbitrary functions $\psi$ for which a certain $\mu$-volume sum converges. Building on this foundation, Einsiedler, Fishman, and Shapira [14], as well as Simmons and Weiss [30], demonstrated that $\mu(W_{d}(\varepsilon\psi_{1/d},\mathbf{0}))$ has full measure for any $\varepsilon>0$, with $\mu$ being either a self-similar measure or an ergodic measure. A complete Khintchine-type theorem for a large class of self-similar measures was established by Khalil and Luethi [20], as well as by Datta and Jana [13], using different techniques. The case of all strongly irreducible self-similar measures was finally settled by Bénard, He and Zhang [3, 2]:
Theorem BHZ ([3, 2]). Let $\mu$ be strongly irreducible self-similar measure. Let $\psi:\mathbb{N}\to\mathbb{R}_{+}$ be a monotonic function. Then,
$\mu\big(W_{d}(\psi,\mathbf{0})\big)=\begin{cases}0&\sum_{q=1}^{\infty}\psi(q)^{d}<\infty,\\[4.30554pt] 1&\sum_{q=1}^{\infty}\psi(q)^{d}=\infty.\end{cases}$
In the inhomogeneous setting, Chow and Yu [12] established that the same zero–one law holds for any $\bm{\theta}\in\mathbb{R}^{d}$, not just for $\bm{\theta}=\bm{0}$, provided that $\dim_{l^{1}}\mu>d-\frac{d-1}{d+1}$. The precise definition of $\dim_{l^{1}}\mu$ will be given in Section 6.2.
For the Hausdorff dimension, comparatively little is known, and the corresponding problems appear to be at least as challenging as those encountered in the context of Khintchine-type theorem for self-similar measures. Levesley, Salp and Velani [26] first established that, for any missing digits set $K$, the Hausdorff dimension of its intersection with $W_{1}(\psi_{\tau},0)$ is at least $(\operatorname{\dim_{H}}K)/(1+\tau)$. In fact, the Cantor-type subset constructed in their proof only guarantees that points are well-approximated by rational points lying within the missing digits set $K$ itself, which are far fewer than rational points in $\mathbb{R}$. Therefore, this lower bound is far from optimal. They further posed the following conjecture, which has recently been fully resolved by Chen [10]. Define the set of very well approximable points by
$\operatorname{VWA}_{d}({\bm{\theta}}):=\bigcup_{\tau>1/d}W_{d}(\psi_{\tau},{\bm{\theta}}).$
Theorem C ([10]). Let $K$ be a self-similar set on $\mathbb{R}$. Then,
$\operatorname{\dim_{H}}\big(K\cap\operatorname{VWA}_{1}(0)\big)=\operatorname{\dim_{H}}K.$
Let $K_{1/3}$ be the middle-third Cantor set. Bugeaud and Durand [9] conjectured that the Hausdorff dimension of $K_{1/3}\cap W_{1}(\psi_{\tau},0)$ should be given by the following explicit formula.
Conjecture BD ([9]). For any $\tau>1$,
$\operatorname{\dim_{H}}\big(K_{1/3}\cap W_{1}(\psi_{\tau},0)\big)=\max\bigg\{\frac{\log 2}{\log 3}+\frac{2}{1+\tau}-1,\ \frac{\log 2/\log 3}{1+\tau}\bigg\}.$
For any $\tau>1$ lying below a certain threshold, Chen [10] provided a sharp upper bound and a comparatively weaker lower bound for the Hausdorff dimension of $K\cap W_{1}(\psi_{\tau},0)$ for all self-similar sets $K$, not only for the middle-third Cantor set. Another major advance was made by Yu [31], who in fact completed this line of research several years before Chen [10] by showing the following theorem. His results apply to sufficiently thick missing digits sets, but do not cover the middle-third Cantor set.
Theorem Y ([31]). Let $K\subset\mathbb{R}$ be a missing digits set, and let $\mu$ be the restriction of $\mathcal{H}^{\operatorname{\dim_{H}}K}$ to $K$. If $\dim_{l^{1}}\mu>1/2$, then
(1.1) $\operatorname{\dim_{H}}\big(K\cap\operatorname{VWA}_{1}(0)\big)=\operatorname{\dim_{H}}K.$
Suppose further that $\operatorname{\dim_{H}}K\dim_{l^{1}}\mu>1/2$, then there exists a constant $\alpha>1$ such that for any $\tau\in(1,\alpha)$,
(1.2) $\operatorname{\dim_{H}}\big(K\cap W_{1}(\psi_{\tau},0)\big)=\operatorname{\dim_{H}}K+\frac{2}{1+\tau}-1.$
Let $\mu_{1/3}$ be the restriction of $\mathcal{H}^{\frac{\log 2}{\log 3}}$ to $K_{1/3}$. It is currently unclear whether $\dim_{l^{1}}\mu_{1/3}>1/2$, and consequently (1.1) in Theorem Y cannot be applied to $K_{1/3}$. Although the result for the Hausdorff dimension of $K_{1/3}\cap\operatorname{VWA}_{1}(0)$ has already been established by Chen [10], Yu’s approach still has its merits, and with suitable refinements, it can be extended to the inhomogeneous setting (see Theorem 1.5 (1)). The second conclusion (1.2) in Theorem Y does not hold for $K_{1/3}$, since $\dim_{l^{1}}\mu_{1/3}\leq\operatorname{\dim_{H}}K_{1/3}$ (see [31, Lemma 1.4 (1) and §3.6]), and hence $\operatorname{\dim_{H}}K_{1/3}\dim_{l^{1}}\mu_{1/3}\leq(\frac{\log 2}{\log 3})^{2}\approx 0.63093^{2}<1/2$. Later, Chow, Varjú and Yu [11] showed that (1.1) and (1.2) in Theorem Y can be respectively applied to 5-ary and 7-ary missing digits sets, each with exactly one missing digit.
The starting point of this paper is inspired by the aforementioned results on Hausdorff dimension. Our goal is to develop a general framework that not only encompasses these previous results but also allows for their natural generalization to higher dimensions, to the inhomogeneous setting, and to results concerning the corresponding Hausdorff measures.
Throughout, the symbols $\ll$ and $\gg$ will be used to indicate an inequality with an unspecified positive multiplicative constant. By $a\asymp b$ we mean $a\ll b$ and $b\ll a$. Let $|E|$ denote the diameter of a set $E$. Let $K\subset\mathbb{R}^{d}$ be a compact subset equipped with a non-atomic probability measure $\mu$. Suppose there exists a constant $\delta>0$ such that, for any $\mathbf{x}\in K$ and any $0<r<|K|$,
$\mu\bigl(B(\mathbf{x},r)\bigr)\asymp r^{\delta},$
where the implied constants are independent of $\mathbf{x}$ and $r$. A measure satisfying this condition is said to be $\delta$-Ahlfors regular.
For any ${\bm{\theta}}\in\mathbb{R}^{d}$, $Q>1$ and $\eta>0$, define
[…]
We now introduce a local counting property which captures the distribution of the sets $A_{Q}(\eta,{\bm{\theta}})$ at small scales with respect to the measure $\mu$.
Definition 1.1.
Let $\alpha>1/d$, $\beta>0$ and ${\bm{\theta}}\in\mathbb{R}^{d}$. We say that the measure $\mu$ satisfies the $(\alpha,\beta,\bm{\theta})$-local counting property if for all sufficiently large $Q\geq 1$ (depending only on $\alpha$ and $\beta$), all $Q^{-\alpha}\leq\eta\leq Q^{-1/d}$, and all balls $B$ whose centres lie in the support of $\mu$ and satisfy $|B|\geq Q^{-\beta}$, one has
(1.3) $\mu\bigl(B\cap A_{Q}(\eta,{\bm{\theta}})\bigr)\asymp\mu(B)\cdot Q\eta^{d},$
where the implied constants are independent of $Q$, $\eta$ and $B$.
Remark 1.
The parameters $\alpha$ and $\beta$ in Definition 1.1 are not independent. To see this, suppose for simplicity that $\mu$ is a $\delta$-Ahlfors regular measure. Applying (1.3) with $\eta=Q^{-\alpha}$, we in particular obtain that $B\cap A_{Q}(Q^{-\alpha},\bm{\theta})\neq\varnothing$. Hence, $B\cap A_{Q}(4Q^{-\alpha},\bm{\theta})$ contains a ball of radius $Q^{-1-\alpha}$. Since $\mu$ is $\delta$-Ahlfors regular, it follows that
$\mu\bigl(B\cap A_{Q}(4Q^{-\alpha},\bm{\theta})\bigr)\gg Q^{-\delta(1+\alpha)}.$
On the other hand, the $(\alpha,\beta,\bm{\theta})$-local counting property yields
$\mu\bigl(B\cap A_{Q}(4Q^{-\alpha},\bm{\theta})\bigr)\asymp\mu(B)\cdot Q^{1-\alpha d}.$
Combining these estimates and using the lower bound $|B|\geq Q^{-\beta}$, together with the $\delta$-Ahlfors regularity of $\mu$, we obtain the necessary condition
$Q^{-\beta\delta}\cdot Q^{1-\alpha d}\gg Q^{-\delta(1+\alpha)}\quad\Longrightarrow\quad\alpha\leq\frac{1+\delta-\beta\delta}{d-\delta}.$
The following theorem provides a sharp upper bound and a reasonably good lower bound for the Hausdorff dimension of well-approximable points on a compact set $K\subset\mathbb{R}^{d}$ whose measure, supported on $K$, satisfies $(\alpha,\beta,{\bm{\theta}})$-local counting property.
Theorem 1.2.
Let $K\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular probability measure $\mu$. Suppose that $\mu$ satisfies the $(\alpha,\beta,\bm{\theta})$-local counting property for some $1/d<\alpha<\frac{1+\delta-\beta\delta}{d-\delta}$, $0<\beta<1$ and ${\bm{\theta}}\in\mathbb{R}^{d}$. Let $\psi$ be a monotonic approximation function. If $1/d<\tau(\psi)<\alpha$, then
(1.4) $\operatorname{\dim_{H}}\big(K\cap W_{d}(\psi,{\bm{\theta}})\big)\leq\delta+\frac{d+1}{1+\tau(\psi)}-d,$
and
(1.5) $\begin{split}&\operatorname{\dim_{H}}\big(K\cap W_{d}(\psi,{\bm{\theta}})\big)\\ \geq&\min\biggl\{\delta-\frac{(\tau(\psi)d-1)(\delta+\frac{1+d}{1+\tau(\psi)}-d)}{\beta\delta},\ \delta+\frac{1+d}{1+\tau(\psi)}-d\biggr\}.\end{split}$
Moreover, if $\delta>d-1$, ${\bm{\theta}}=\mathbf{0}$ and $\psi(q)\geq cq^{-\tau}$ with $1/d<\tau<\min\{\alpha,\frac{\delta+2-d}{2d-\delta-1}\}$ for infinitely many $q$. Then,
(1.6) $\mathcal{H}^{s}\big(K\cap W_{d}(\psi,\bf 0)\big)=\infty,$
where
(1.7) $s=\min\bigg\{\delta-\frac{(\tau-\frac{1}{d})(2d-\delta-1)(\delta+\frac{1+d}{1+\tau}-d)}{\beta\delta},\ \delta+\frac{1+d}{1+\tau}-d\bigg\},$
is hence a lower bound for $\operatorname{\dim_{H}}(K\cap W_{d}(\psi,\bf 0))$.
Remark 2.
By Conjecture BD, the upper bound (1.4) is sharp. The two lower bounds described in (1.5) and (1.7), on the other hand, are relatively weaker. In fact, when $\delta>d-1$, the lower bound described in (1.7) is strictly larger than that in (1.5) since
$(\tau-1/d)(2d-\delta-1)<(\tau-1/d)\cdot d=\tau d-1,$
and in certain cases, this bound, namely (1.7), is even sharp. Specifically, as we will see in Remark 8, the minimum in (1.7) is $\delta+\frac{1+d}{1+\tau}-d$ provided that
$\delta\geq 2d-1-\frac{d^{2}\beta}{1+d}\mathrm{\quad and\quad}\beta>1-d^{-2}.$
It is easy to observe that the most restrictive condition here is on $\beta$. When $d=1$, we only require $\beta>0$, which is relatively easy to satisfy. However, as $d$ increases, $\beta$ must also increase and approach $1$. It is unclear, even for the Lebesgue measure, whether the corresponding $\beta$ can be taken sufficiently close to $1$ when $d>1$. A crucial reason why the lower bound is more effective in one dimension is the simple fact that the distance between any two distinct rationals $p/q$ and $p^{\prime}/q^{\prime}$ is at least $1/(qq^{\prime})$. Although there are analogous results (see Lemma 4.1) in higher dimensions, they are not strong enough, so it is necessary to require $\beta>1-d^{-2}$.
Remark 3.
It is easy to see that, in Theorem 1.2, the larger the value of $\beta$, the better the lower bounds. However, in general, $\beta$ tends to be quite small. If one hopes to apply this result to Conjecture BD, it would require $\beta>2(1-\frac{\log 2}{\log 3})\approx 0.7381$. According to the currently known results, achieving such a large value of $\beta$ seems to be very difficult, see also Remark 6.
Remark 4.
The appearance of $\beta$ in the two lower bounds (1.5) and (1.7) stems from the following reason. When analyzing the fractal structure of the set $K\cap W_{d}(\psi,{\bm{\theta}})$, the $(\alpha,\beta,{\bm{\theta}})$-local counting property allows us to obtain good estimates for balls with radius larger than $Q^{-\beta}$. However, for balls with smaller radius, such estimates are no longer available, and one must resort to worst-case estimates. This limitation is what can cause the lower bound to be smaller than the value one might ideally expect.
Next, we present several consequences, which, as we shall see later, extend the works of Chen [10] and Yu [31]. Note that when $d=1$, the additional restriction on $\tau$ (imposed so as to obtain the better lower bound (1.7)) is
$1<\tau<\min\left\{\alpha,\ \frac{1+\delta}{1-\delta}\right\}=\alpha,$
since $\alpha<\frac{1+\delta-\beta\delta}{1-\delta}$.
Corollary 1.3.
Let $K\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular probability measure $\mu$. Suppose that $\mu$ satisfies the $(\alpha,\beta,\bm{\theta})$-local counting property for some $1/d<\alpha<\frac{1+\delta-\beta\delta}{d-\delta}$, $0<\beta<1$ and ${\bm{\theta}}\in\mathbb{R}^{d}$. Then, the following statements hold.
  1. (1)
    $\operatorname{\dim_{H}}\big(K\cap\operatorname{VWA}_{d}({\bm{\theta}})\big)=\operatorname{\dim_{H}}K=\delta.$
  2. (2)
    If $d=1$, $\theta=0$, $\beta>0$ and $\delta>1-\beta/2$, then for any $1<\tau<\alpha$,
    $\mathcal{H}^{\delta+\frac{2}{1+\tau}-1}\big(K\cap E_{1}(\tau,0)\big)=\infty.$
Based on this corollary, we provide several concrete examples.
Theorem 1.4.
Let $K\subset\mathbb{R}^{d}$ be a strongly irreducible self-similar set satisfying the open set condition. Suppose that ${\bm{\theta}}=\mathbf{0}$. Then, the following statements hold.
  1. (1)
    For any $d\geq 1$, we have $\operatorname{\dim_{H}}(K\cap\operatorname{VWA}_{d}(\mathbf{0}))=\operatorname{\dim_{H}}K$.
  2. (2)
    Let $d=1$. There exist computable constants $0<\delta_{0}<1$ and $\alpha>1$ such that for any $K$ with $\operatorname{\dim_{H}}K\geq\delta_{0}$, we have
    $\mathcal{H}^{\delta+\frac{2}{1+\tau}-1}\big(K\cap E_{1}(\tau,0)\big)=\infty,\quad\text{for any $\tau\in(1,\alpha)$.}$
Remark 5.
Theorem 1.4 (1) extends Theorem C to higher dimensions. Moreover, for sufficiently thick self-similar sets, we are able to obtain the Hausdorff measure of the set $K\cap E_{1}(\tau,0)$, which generalizes Theorem Y on missing digits sets to thick self-similar sets. However, Theorem 1.4 (2) cannot be straightforwardly extended to higher dimensions, as explained in Remarks 2 and 8.
Remark 6.
The constants $\delta_{0}$ and $\alpha$ in Theorem 1.4 (2) are, in principle, computable. They depend on the exponent $\kappa$ appearing in the error term of the effective equidistribution result established by Bénard, He and Zhang [3, 2] (see Theorem 6.1). Roughly speaking, the admissible range of the parameter $\beta$ depends on the exponent $\kappa$, and increases as $\kappa$ becomes larger. More precisely, by Corollary 6.5, $\beta$ can be taken arbitrarily close to $\frac{2\kappa}{2+\kappa}$ from below. However, the authors of [3, 2] do not provide an explicit expression for $\kappa$, and in fact the available value of $\kappa$ is very small. In order to apply the result to the middle-third Cantor set, it follows from Corollary 1.3 (2) that one would require $\delta=\frac{\log 2}{\log 3}>1-\beta/2>1-\frac{\kappa}{2+\kappa}$, or equivalently $\kappa>2(1-\frac{\log 2}{\log 3})/\frac{\log 2}{\log 3}\approx 1.1699$, which seems difficult to achieve using their method. For this reason, we do not attempt to compute the explicit values of $\delta_{0}$ and $\alpha$ here.
Theorem 1.5.
Let $K\subset\mathbb{R}$ be a missing digits set, and let $\mu$ be the restriction of $\mathcal{H}^{\operatorname{\dim_{H}}K}$ to $K$. Then, the following statements hold.
  1. (1)
    If $\dim_{l^{1}}\mu>1/2$, then for any $\theta\in\mathbb{R}$,
    $\operatorname{\dim_{H}}\big(K\cap\operatorname{VWA}_{1}(\theta)\big)=\operatorname{\dim_{H}}K.$
  2. (2)
    If $\operatorname{\dim_{H}}K\dim_{l^{1}}\mu>1/2$, then there exists $\alpha>1$ such that
    $\mathcal{H}^{\delta+\frac{2}{1+\tau}-1}\big(K\cap E_{1}(\tau,0)\big)=\infty,\quad\text{for any $\tau\in(1,\alpha)$.}$
Remark 7.
Item (1) of Theorem 1.5 provides an affirmative answer to a question posed by Yu [31, Remark 9.4] concerning inhomogeneous Diophantine approximation on missing digits sets. Item (2) goes further by strengthening the conclusion of Theorem Y from Hausdorff dimension to Hausdorff measure. As mentioned earlier, it follows from the work of Chow, Varjú, and Yu [11, Propositions 2.4 and 2.5] that items (1) and (2) apply respectively to 5-ary and 7-ary missing digits sets, each with exactly one missing digit.
Organization of the paper
In Section 2, we introduce a useful tool for proving Theorem 1.2. Sections 3 and 4 are devoted to proving Theorem 1.2, while Section 5 presents the proof of Corollary 1.3. In Section 6, we focus on two specific classes of sets: self-similar sets and missing digits sets, which are treated in two separate subsections. More precisely, we verify that these two classes satisfy the $(\alpha,\beta,{\bm{\theta}})$-local counting property for some parameters $\alpha$, $\beta$ and ${\bm{\theta}}\in\mathbb{R}^{d}$, respectively, which in turn yields Theorems 1.4 and 1.5.
2. Hausdorff measure and content
Here and hereafter, we will assume that $K\subset\mathbb{R}^{d}$ is a compact subset equipped with a $\delta$-Ahlfors regular probability measure $\mu$.
In this section, we introduce a useful tool for estimating the Hausdorff dimension and measure of the set $K\cap W_{d}(\psi,{\bm{\theta}})$.
Let $0<s\leq d$. For a set $E\subset\mathbb{R}^{d}$ and $\eta>0$, let
$\mathcal{H}_{\eta}^{s}(E):=\inf\bigg\{\sum_{i}|B_{i}|^{s}:E\subset\bigcup_{i\geq 1}B_{i},\text{ where $B_{i}$ are balls with $|B_{i}|\leq\eta$}\bigg\}.$
The $s$-dimensional Hausdorff measure of $E$ is defined as
$\mathcal{H}^{s}(E):=\lim_{\eta\to 0^{+}}\mathcal{H}_{\eta}^{s}(E).$
When $\eta=\infty$, $\mathcal{H}_{\infty}^{s}(E)$ is referred to as $s$-dimensional Hausdorff content of $E$.
In [15], Falconer introduced and systematically developed the notion of the large intersection property, which provides a powerful theoretical framework for studying the Hausdorff dimension of $\limsup$ sets. This property ensures that certain subsets maintain large Hausdorff dimension even under countable intersections, making it particularly useful in Diophantine approximation.
Definition 2.1 ([15]).
Let $0<s\leq\operatorname{\dim_{H}}K$. We define $\mathscr{G}_{\dim}^{s}(K)$ to be the class of $G_{\delta}$-subsets $E$ of $K$ such that there exists a constant $c>0$ satisfying the following property: for any $0<t<s$ and any ball $B\subset K$,
(2.1) $\mathcal{H}_{\infty}^{t}(E\cap B)>c\mathcal{H}_{\infty}^{t}(B).$
Intuitively, the condition (2.1) ensures that the set $E$ is “large” in every ball, in the sense that it occupies a uniformly positive proportion of the $t$-dimensional Hausdorff content of $B$ for all $t<s$.
Theorem 2.2 ([15]).
Let $0<s\leq d$. The class $\mathscr{G}_{\dim}^{s}(K)$ is closed under countable intersections. Moreover, for any $E\in\mathscr{G}_{\dim}^{s}(K)$, we have
$\operatorname{\dim_{H}}E\geq s.$
However, Falconer’s original definition could only provide lower bounds for the Hausdorff dimension, not the Hausdorff measure. This limitation was recently overcome by the first author in [16], who introduced a refined version of Definition 2.1.
Definition 2.3 ([16, Definition 2.3]).
Let $0<s\leq\operatorname{\dim_{H}}K$. We define $\mathscr{G}_{\mathrm{mea}}^{s}(K)$ to be the class of $G_{\delta}$-subsets $E$ of $K$ such that there exists a constant $c>0$ satisfying the following property: for any ball $B\subset K$,
(2.2) $\mathcal{H}_{\infty}^{s}(E\cap B)>c\mathcal{H}_{\infty}^{s}(B).$
The difference lies in whether the critical value $t=s$ can be attained. If the critical value cannot be attained, it is therefore natural that one cannot obtain information about the Hausdorff measure, as the Hausdorff measure is particularly sensitive to the critical dimension. On the other hand, if the critical value can be attained, one may expect to obtain information about the Hausdorff measure, as described in the theorem below.
Theorem 2.4 ([16, Theorem 2.4]).
Let $0<s\leq d$. The class $\mathscr{G}_{\mathrm{mea}}^{s}(K)$ is closed under countable intersections. Moreover, for any $E\in\mathscr{G}_{\mathrm{mea}}^{s}(K)$, we have
$\mathcal{H}^{s}(E)=\mathcal{H}^{s}(K).$
Inspired by the mass transference principle from balls to open sets [24, 33], the author [16] derived conditions that are significantly weaker than the original definitions, while still yielding the corresponding large intersection property.
Theorem 2.5 ([16, Corollary 2.6]).
Let $0<s\leq d$. Assume that $\{B_{k}\}$ is a sequence of balls in $K$ with radii tending to 0, and that $\mu(\limsup B_{k})=1$. Let $\{E_{n}\}$ be a sequence of open sets (not necessarily contained in $B_{k}$). The following statement hold.
  1. (1)
    If for any $0<t<s$, there exists a constant $c_{t}>0$ such that for any $B_{k}$,
    $\limsup_{n\to\infty}\mathcal{H}_{\infty}^{t}(E_{n}\cap B_{k})>c_{t}\mu(B_{k}),$
    then $\limsup E_{n}\in\mathscr{G}_{\mathrm{dim}}^{s}(K)$.
  2. (2)
    If there exists a constant $c_{s}>0$ such that for any $B_{k}$,
    $\limsup_{n\to\infty}\mathcal{H}_{\infty}^{s}(E_{n}\cap B_{k})>c_{s}\mu(B_{k}),$
    then $\limsup E_{n}\in\mathscr{G}_{\mathrm{mea}}^{s}(K)$.
With this result now at our disposal, the main ideas to prove Theorem 1.2 are to verify certain sets under consideration satisfying some Hausdorff content bound. For this purpose, the following mass distribution principle will be crucial.
Proposition 2.6 (Mass distribution principle [6, Lemma 1.2.8]).
Let $E$ be a Borel subset of $\mathbb{R}^{d}$. If $E$ supports a Borel probability measure $\nu$ that satisfies
$\nu(B)\leq c|B|^{s},$
for some constant $0<c<\infty$ and for every ball $B$, then $\mathcal{H}_{\infty}^{s}(E)\geq 1/c$.
3. Proof of Theorem 1.2: upper bound of $\operatorname{\dim_{H}}\big(K\cap W_{d}(\psi,\theta)\big)$
Before giving the proof of the upper bound for $\operatorname{\dim_{H}}(K\cap W_{d}(\psi,{\bm{\theta}}))$, we first investigate the geometric structure of the set $K\cap A_{Q}(\eta,\bm{\theta})$. In particular, we show that, under a counting property weaker than the $(\alpha,\beta,\bm{\theta})$-local counting property, this set admits sharp covering estimates.
Lemma 3.1.
Suppose that for all sufficiently large $Q\geq 1$ and all $Q^{-\alpha}\leq\eta\leq Q^{-1/d}$, one has
(3.1) $\mu\bigl(A_{Q}(\eta,{\bm{\theta}})\bigr)\asymp Q\eta^{d},$
where the implied constants are independent of $Q$ and $\eta$. Then, for all sufficiently large $Q\geq 1$ and all $Q^{-\alpha}\leq\eta\leq Q^{-1/d}$, the set $K\cap A_{Q}(\eta,{\bm{\theta}})$ can be covered by
$\asymp Q^{1+\delta}\eta^{d-\delta}$
balls of radius $Q^{-1}\eta$.
Proof.
The conclusion does not follow directly from (3.1), since the ‘rational’ points $(\mathbf{p}+{\bm{\theta}})/q$ need not lie in $K$, and therefore the size of the intersection
$K\cap B\left(\frac{\mathbf{p}+{\bm{\theta}}}{q},\,\frac{\eta}{q}\right)$
may vary significantly for different choices of $(\mathbf{p}+{\bm{\theta}})/q$.
Let $\mathbf{x}\in K\cap A_{Q}(\eta,{\bm{\theta}})$. By definition, there exists $(\mathbf{p}+{\bm{\theta}})/q$ with $Q\leq q<2Q$ such that
$\left|\mathbf{x}-\frac{\mathbf{p}+{\bm{\theta}}}{q}\right|<\frac{\eta}{q}\leq\frac{\eta}{Q}.$
If $\mathbf{y}$ satisfies $|\mathbf{x}-\mathbf{y}|<Q^{-1}\eta$, then
$\left|\mathbf{y}-\frac{\mathbf{p}+{\bm{\theta}}}{q}\right|\leq|\mathbf{y}-\mathbf{x}|+\left|\mathbf{x}-\frac{\mathbf{p}+{\bm{\theta}}}{q}\right|<\frac{2\eta}{Q}<\frac{4\eta}{q},$
which implies that $\mathbf{y}\in A_{Q}(4\eta,{\bm{\theta}})$. Consequently,
(3.2) $K\cap A_{Q}(\eta,{\bm{\theta}})\subset\bigcup_{\mathbf{x}\in K\cap A_{Q}(\eta,{\bm{\theta}})}B(\mathbf{x},Q^{-1}\eta)\subset K\cap A_{Q}(4\eta,{\bm{\theta}}).$
By our assumption (3.1), the $\mu$-measures of the two outer sets in (3.2) are comparable. Therefore,
(3.3) $\mu\bigg(\bigcup_{\mathbf{x}\in K\cap A_{Q}(\eta,{\bm{\theta}})}B(\mathbf{x},Q^{-1}\eta)\bigg)\asymp Q\eta^{d}.$
By the $\delta$-Ahlfors regularity of $\mu$, this implies that $K\cap A_{Q}(\eta,{\bm{\theta}})$ can be covered by
$\asymp\frac{Q\eta^{d}}{(Q^{-1}\eta)^{\delta}}=Q^{1+\delta}\eta^{d-\delta}$
balls of radius $Q^{-1}\eta$. ∎
Let $\alpha$ be as in Theorem 1.2, and assume that $1/d<\tau(\psi)<\alpha$. Obviously, the $(\alpha,\beta,\bm{\theta})$-local counting property implies the global estimate (3.1), and hence Lemma 3.1 provides a sharp covering of $K\cap A_{Q}(\eta,\bm{\theta})$ with $Q^{-\alpha}\leq\eta\leq Q^{-1/d}$. Accordingly, to obtain the upper bound of $K\cap W_{d}(\psi,{\bm{\theta}})$, we define
$\phi(q)=\max\{\psi(q),q^{-\alpha}\}.$
It is evident that $K\cap W_{d}(\psi,{\bm{\theta}})\subset K\cap W_{d}(\phi,{\bm{\theta}})$ and $\phi$ inherits the monotonicity from $\psi$. Moreover, since $\tau(\psi)<\alpha$, we have
$\tau(\phi)=\tau(\psi)<\alpha.$
In what follows, we bound the Hausdorff dimension of $K\cap W_{d}(\phi,{\bm{\theta}})$ from above, which immediately gives an upper bound for that of $K\cap W_{d}(\psi,{\bm{\theta}})$.
By the monotonicity of $\phi$, it is straightforward to verify that if $\mathbf{x}\in K\cap W_{d}(\phi,{\bm{\theta}})$, then there exist infinitely many integers $m$ such that
$\|q\mathbf{x}-{\bm{\theta}}\|<\phi(2^{m})\quad\text{for some $2^{m}\leq q<2^{m+1}$}.$
Equivalently, $\mathbf{x}\in K\cap A_{2^{m}}(\phi(2^{m}),{\bm{\theta}})$ for infinitely many $m$. Therefore, we have
$K\cap W_{d}(\phi,{\bm{\theta}})\subset\bigcap_{M=1}^{\infty}\bigcup_{m=M}^{\infty}K\cap A_{2^{m}}\big(\phi(2^{m}),{\bm{\theta}}\big).$
Note that for each $M\geq 1$, the set $\bigcup_{m=M}^{\infty}K\cap A_{2^{m}}(\phi(2^{m}),{\bm{\theta}})$ is a cover of $K\cap W_{d}(\phi,{\bm{\theta}})$. Since $\phi(2^{m})\geq 2^{-m\alpha}$ for all $m\geq 1$, Lemma 3.1 is applicable to $K\cap A_{2^{m}}(\phi(2^{m}),{\bm{\theta}})$ whenever $m$ is sufficiently large. Therefore, for all sufficiently large $m$, it follows from Lemma 3.1 that $K\cap A_{2^{m}}(\phi(2^{m}),{\bm{\theta}})$ can be covered by
$\asymp 2^{m(1+\delta)}\phi(2^{m})^{d-\delta}$
balls of radius $2^{-m}\phi(2^{m})$. Let $1/d<\tau<\tau(\phi)$. Then, for all sufficiently large $q$, we have $\phi(q)\leq q^{-\tau}$. For any $0<s\leq\delta$, by the definition of $s$-dimensional Hausdorff measure,
$\begin{split}\mathcal{H}^{s}\big(K\cap W_{d}(\phi,{\bm{\theta}})\big)&\ll\liminf_{M\to\infty}\sum_{m=M}^{\infty}2^{m(1+\delta)}\phi(2^{m})^{d-\delta}\cdot\big(2^{-m}\phi(2^{m})\big)^{s}\\ &\leq\liminf_{M\to\infty}\sum_{m=M}^{\infty}2^{m(1+\delta)}2^{-m\tau(d-\delta)}\cdot(2^{-m}2^{-m\tau})^{s}\\ &=\liminf_{M\to\infty}\sum_{m=M}^{\infty}2^{m(1+\delta-\tau(d-\delta)-s(1+\tau))}.\end{split}$
Observe that the above $\liminf$ tends to zero if and only if the exponent satisfies
$1+\delta-\tau(d-\delta)-s(1+\tau)<0,$
which is equivalent to
$s>\frac{-\tau(d-\delta)+1+\delta}{1+\tau}=\delta+\frac{1+d}{1+\tau}-d.$
Therefore,
$\operatorname{\dim_{H}}\big(K\cap W_{d}(\phi,{\bm{\theta}})\big)\leq\delta+\frac{1+d}{1+\tau}-d.$
Since $\tau<\tau(\phi)=\tau(\psi)$ is arbitrary, we conclude that
$\operatorname{\dim_{H}}\big(K\cap W_{d}(\phi,{\bm{\theta}})\big)\leq\delta+\frac{1+d}{1+\tau(\phi)}-d,$
which completes the proof of (1.4) in Theorem 1.2.
4. Proof of Theorem 1.2: lower bound of $\operatorname{\dim_{H}}\big(K\cap W_{d}(\psi,\theta)\big)$
The goal of this section is to prove the lower bound described in Theorem 1.2. By the monotonicity of $\psi$, it is straightforward to verify that if $\mathbf{x}\in K$ and there exist infinitely many integers $m$ such that
$\|q\mathbf{x}-{\bm{\theta}}\|<\psi(2^{m+1})\quad\text{for some $2^{m}\leq q<2^{m+1}$},$
then $\mathbf{x}\in K\cap W_{d}(\psi,{\bm{\theta}})$. Therefore, we have
(4.1) $\begin{split}K\cap W_{d}(\psi,{\bm{\theta}})&\supset\bigcap_{M=1}^{\infty}\bigcup_{m=M}^{\infty}K\cap A_{2^{m}}\big(\psi(2^{m+1}),{\bm{\theta}}\big)\\ &=\limsup_{m\to\infty}K\cap A_{2^{m}}\big(\psi(2^{m+1}),{\bm{\theta}}\big).\end{split}$
To establish a lower bound for the Hausdorff dimension (or measure) of $K\cap W_{d}(\psi,{\bm{\theta}})$, it suffices, by Theorem 2.5, to show that for some suitable parameter $s<\delta$,
$\limsup_{m\to\infty}\mathcal{H}_{\infty}^{s}\Big(B\cap A_{2^{m}}\big(\psi(2^{m+1}),{\bm{\theta}}\big)\Big)\gg\mu(B)$
holds for all ball $B$ in $K$, where the implied constant is independent of $B$.
Now, assume that $\mu$ satisfies the $(\alpha,\beta,\bm{\theta})$-local counting property for some $1/d<\alpha<\frac{1+\delta-\beta\delta}{d-\delta}$, $0<\beta<1$ and ${\bm{\theta}}\in\mathbb{R}^{d}$.
Let $B\subset K$ be fixed. Choose $\tau$ so that $1/d<\tau(\psi)<\tau<\alpha$, this is possible due to our assumption $\tau(\psi)<\alpha$. Hence, for any $c>0$, there exists infinitely many $m$ such that
(4.2) $\psi(2^{m+1})>c2^{-m\tau}>2^{-m\alpha}.$
Here we have introduced the constant $c>0$ so that, after completing the proof of (1.5) in Section 4.1, the same construction can be used with minimal modification to establish (1.7) in Section 4.2.
Let $m\geq\log_{2}|B|^{-1/\beta}$ be an arbitrary integer satisfying (4.2) and such that the local estimate (1.3) can be applied with $Q$ replaced by $2^{m}$. Note that there are still infinitely many admissible choices of $m$. For simplicity, write
(4.3) $Q=2^{m}\mathrm{\quad and\quad}F=B\cap A_{Q}(cQ^{-\tau},{\bm{\theta}}).$
Define a probability measure supported on $F\subset B\cap A_{2^{m}}(\psi(2^{m+1}),{\bm{\theta}})$ by
(4.4) $\nu:=\frac{\mu|_{F}}{\mu(F)}.$
Since $m\geq\log_{2}|B|^{-1/\beta}$ (equivalently $|B|\geq 2^{-m\beta}=Q^{-\beta}$) and $cQ^{-\tau}>Q^{-\alpha}$ (see (4.2)), by the $(\alpha,\beta,{\bm{\theta}})$-local counting property of $\mu$,
(4.5) $\mu(F)=\mu\bigl(B\cap A_{Q}(cQ^{-\tau},{\bm{\theta}})\bigr)\asymp\mu(B)\cdot Q^{1-\tau d}.$
Next, we estimate the $\nu$-measure of an arbitrary ball. Let $\mathbf{x}\in F$ and let $0<r<|B|$. We proceed by considering several ranges of $r>0$.
Case 1: $Q^{-\beta}\leq r<|B|$. In this case, the $(\alpha,\beta,{\bm{\theta}})$-local counting property is applicable to the ball $B(\mathbf{x},r)$. By (1.3) and the $\delta$-Ahlfors regularity of $\mu$, we have
(4.6) $\mu\big(B(\mathbf{x},r)\cap A_{Q}(cQ^{-\tau},{\bm{\theta}})\big)\asymp\mu\big(B(\mathbf{x},r)\big)\cdot Q^{1-\tau d}\asymp r^{\delta}\cdot Q^{1-\tau d}.$
Therefore,
[…]
since $\tau>1/d$.
Case 2: $Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}\leq r<Q^{-\beta}$. Then, applying the estimate from Case 1 with $r=Q^{-\beta}$, we obtain
$\nu\big(B(\mathbf{x},r)\big)\leq\nu\big(B(\mathbf{x},Q^{-\beta})\big)\ll\frac{Q^{-\beta\delta}}{\mu(B)}\leq\frac{r^{\delta+\frac{1+d}{1+\tau}-d}}{\mu(B)}.$
The $(\alpha,\beta,{\bm{\theta}})$-local counting property allows us to control the $\nu$-measure of a ball $B(\mathbf{x},r)$ effectively when $r\geq Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}$. However, this estimate breaks down at smaller scales. To deal with this difficulty, we will adopt different approaches depending on whether the conditions $\delta>d-1$, ${\bm{\theta}}=\mathbf{0}$ and $\tau<\frac{\delta+2-d}{2d-\delta-1}$ are satisfied. Accordingly, the analysis will be divided into separate cases in the following two subsections.
4.1. Estimate of the $\nu$-measure of balls for general cases
Case 3A: $0<r<Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}$. At this scale, the estimate (4.6) is not guaranteed to hold, and we instead have the following somewhat coarse estimate:
[…]
Completing the proof of (1.5) in Theorem 1.2.
Let
$s_{\tau}=\min\biggl\{\delta-\frac{(\tau d-1)(\delta+\frac{1+d}{1+\tau}-d)}{\beta\delta},\ \delta+\frac{1+d}{1+\tau}-d\biggr\}.$
By Cases 1, 2 and 3A above, we have
$\nu\big(B(\mathbf{x},r)\big)\ll\frac{r^{s_{\tau}}}{\mu(B)}.$
Note that $\nu$ is supported on $F=B\cap A_{Q}(cQ^{-\tau},{\bm{\theta}})\subset B\cap A_{2^{m}}(\psi(2^{m+1}),{\bm{\theta}})$. Therefore, by mass distribution principle (see Proposition 2.6),
$\mathcal{H}_{\infty}^{s_{\tau}}\big(B\cap A_{2^{m}}(\psi(2^{m+1}),{\bm{\theta}})\big)\gg\mu(B).$
Since there are infinitely many admissible choices of $m$ (see the choice of $m$ at the beginning of this section), it follows that
$\limsup_{m\to\infty}\mathcal{H}_{\infty}^{s_{\tau}}\big(B\cap A_{2^{m}}(\psi(2^{m+1}),{\bm{\theta}})\big)\gg\mu(B)$
holds for any ball $B\subset K$. Let $\tau\downarrow\tau(\psi)$ and denote by $s$ the limit of $s_{\tau}$. Theorem 2.5 (1) together with (4.1) yields
$K\cap W_{d}(\psi,{\bm{\theta}})\in\mathscr{G}_{\mathrm{dim}}^{s}(K),$
and hence
$\begin{split}&\operatorname{\dim_{H}}\big(K\cap W_{d}(\psi,{\bm{\theta}})\big)\\ \geq&s=\min\biggl\{\delta-\frac{(\tau(\psi)d-1)(\delta+\frac{1+d}{1+\tau(\psi)}-d)}{\beta\delta},\ \delta+\frac{1+d}{1+\tau(\psi)}-d\biggr\}.\qed\end{split}$
4.2. Estimate of the $\nu$-measure of balls for $\delta\geq d-1$, $\bf\theta=0$, and $\tau<\frac{\delta+2-d}{2d-\delta-1}$
Parallel to Case 3A, we now continue the discussion from Case 2, dealing with the situation
$0<r<Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}$
by a method different from that in Section 4.1. We divide the range of $r$ into three cases, which will be analyzed separately in Cases 3B–5B below. Prior to discussing these cases, we first state several results that will be used in the subsequent analysis.
Lemma 4.1 ([25, Lemma 4]).
Let $E\subset\mathbb{R}^{d}$ be a convex set with $d$-dimensional Lebesgue measure
$\mathcal{L}_{d}(E)\leq Q^{-(1+d)}/(d!).$
Then the rational points in $E$ with denominators $1\leq q\leq Q$ all lie on some hyperplane in $\mathbb{R}^{d}$.
This lemma can be viewed as a higher-dimensional analogue of the classical one-dimensional fact: any interval in $\mathbb{R}$ of length $\leq Q^{-2}$ contains at most one rational number $p/q$ with $1\leq q\leq Q$.
In [23, Proposition 6.3], Kleinbock and Weiss proved that any $\delta$-Ahlfors regular measure with $\delta>d-1$ is absolutely $(C,\delta+1-d)$-decaying for some constant $C$ depending only on $d$ and the measure $\mu$, in the following sense: for any ball $B(\mathbf{x},r)\subset\mathbb{R}^{d}$, any hyperplane $L\subset\mathbb{R}^{d}$, and any $\varepsilon>0$,
(4.7) $\mu\bigl(B(\mathbf{x},r)\cap L^{(\varepsilon)}\bigr)\leq C\biggl(\frac{\varepsilon}{r}\biggr)^{\delta+1-d}\mu\bigl(B(\mathbf{x},r)\bigr),$
where $L^{(\varepsilon)}$ denotes the $\varepsilon$-neighbourhood of $L$.
Case 3B: $Q^{-(1+1/d)}/(8d!)\leq r<Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}$. Divide the ball $B(\mathbf{x},r)$ into smaller hypercubes of equal side length $Q^{-(1+1/d)}/(8d!)$. If any such hypercube $\mathcal{C}$ has a nonempty intersection with $K$, then $2\mathcal{C}\cap K$ contains a ball of radius $|\mathcal{C}|$, where $2\mathcal{C}$ denotes the hypercube with the same center as $\mathcal{C}$ but twice the side length. Since any point in $K$ belongs to at most $3^{d}$ different sets of the form $2\mathcal{C}\cap K$, a simple volume argument shows that the ball $B(\mathbf{x},r)$ can intersect at most
$\ll\frac{r^{\delta}}{Q^{-(1+1/d)\delta}}$
hypercubes $\mathcal{C}$. Furthermore, since each $\mathcal{C}$ is convex and its $d$-dimensional Lebesgue measure is bounded by $Q^{-(1+d)}/(8d!)^{d}\leq(2Q)^{-(1+d)}/(d!)$, it follows from the simplex lemma above (see Lemma 4.1) that the rational points with denominator $1\leq q\leq 2Q$ contained in $\mathcal{C}$ lie in some hyperplane $L\subset\mathbb{R}^{d}$. Therefore,
(4.8) $\mathcal{C}\cap A_{Q}(cQ^{-\tau},\mathbf{0})\subset\mathcal{C}\cap L^{(cQ^{-(1+\tau)})}.$
Using the absolutely decaying propertying of $\mu$ (see (4.7)), we have
$\begin{split}\mu|_{F}\big(B(\mathbf{x},r)\big)&\leq\mu\big(B(\mathbf{x},r)\cap A_{Q}(cQ^{-\tau},\mathbf{0})\big)\\ &\ll\frac{r^{\delta}}{Q^{-(1+1/d)\delta}}\cdot\bigg(\frac{Q^{-(1+\tau)}}{Q^{-(1+1/d)}}\bigg)^{\delta+1-d}Q^{-(1+1/d)\delta}\\ &=r^{\delta}Q^{(-\tau+1/d)(\delta+1-d)}.\end{split}$
It then follows from the definition of $\nu$ that
[…]
Case 4B: $cQ^{-(1+\tau)}\leq r<Q^{-(1+1/d)}/(8d!)$. Since the $d$-dimensional Lebesgue measure of the ball $B(\mathbf{x},r)$ is less than $Q^{-(1+d)}/(4d!)^{d}$, by the same reason as (4.8),
$B(\mathbf{x},r)\cap A_{Q}(cQ^{-\tau},\mathbf{0})\subset B(\mathbf{x},r)\cap L^{(cQ^{-(1+\tau)})}$
for some hyperplane $L\subset\mathbb{R}^{d}$. Again, by the absolutely decaying property of $\mu$ and the definition of $\nu$, we have
[…]
Since $\tau<\frac{\delta+2-d}{2d-\delta-1}$, the exponent of $Q$ satisfies
$\begin{split}-(1+\tau)(\delta+1-d)-(1-\tau d)=\tau(2d-\delta-1)-(\delta+2-d)<0.\end{split}$
By $cQ^{-(1+\tau)}\leq r$, we have
$Q^{-(1+\tau)(\delta+1-d)-(1-\tau d)}\ll r^{\delta+1-d+\frac{1-\tau d}{1+\tau}}.$
It then follows that
$\nu\big(B(\mathbf{x},r)\big)\ll\frac{r^{\delta+\frac{1+d}{1+\tau}-d}}{\mu(B)}.$
Case 5B: $0<r<cQ^{-(1+\tau)}$. Then,
[…]
Remark 8.
To ensure that the exponent of $r$ in Case 3B satisfies
$\delta-\frac{(\tau-1/d)(2d-\delta-1)\bigl(\delta+\tfrac{1+d}{1+\tau}-d\bigr)}{\beta\delta}\geq\delta+\frac{1+d}{1+\tau}-d,$
it suffices to require that
$\tau\geq\frac{d\beta\delta-(2d-\delta-1)(1+\delta)}{(2d-\delta-1)(\delta-d)}.$
This condition is automatically satisfied if
$\frac{d\beta\delta-(2d-\delta-1)(1+\delta)}{(2d-\delta-1)(\delta-d)}\leq\frac{1}{d}.$
A straightforward simplification of this inequality yields
$\delta\geq 2d-1-\frac{d^{2}\beta}{1+d}.$
Since $\delta<d$, this condition is non-vacuous provided that
$\beta>1-d^{-2}.$
When $d=1$, there are indeed many measures satisfying this condition, since in this case $1-d^{-2}=0$. However, when $d>1$, even in the case of the Lebesgue measure, it remains unclear whether it satisfies this condition. Nevertheless, the above argument still provides some insight into how the higher-dimensional case might be approached.
Completing the proof of (1.6) in Theorem 1.2.
Let
$s=\min\bigg\{\delta-\frac{(\tau-\frac{1}{d})(2d-\delta-1)(\delta+\frac{1+d}{1+\tau}-d)}{\beta\delta},\ \delta+\frac{1+d}{1+\tau}-d\bigg\}.$
By Cases 1, 2, 3B, 4B and 5B above, we have
$\nu\big(B(\mathbf{x},r)\big)\ll\frac{r^{s}}{\mu(B)}.$
Note that $\nu$ is supported on $B\cap A_{Q}(cQ^{-\tau},\mathbf{0})$. Therefore,
$\mathcal{H}_{\infty}^{s}\big(B\cap A_{Q}(cQ^{-\tau},\mathbf{0})\big)\gg\mu(B).$
Since $\psi(q)\geq cq^{-\tau}$ with $1/d<\tau<\min\{\alpha,\frac{\delta+2-d}{2d-\delta-1}\}$ for infinitely many $q$, it follows that
$\limsup_{\begin{subarray}{c}Q\to\infty\\ \psi(Q)\geq cQ^{-\tau}\end{subarray}}\mathcal{H}_{\infty}^{s}\big(B\cap A_{Q}(cQ^{-\tau},\mathbf{0})\big)\gg\mu(B)$
holds for any ball $B\subset K$. By Theorem 2.5 (2),
$\limsup_{\begin{subarray}{c}Q\to\infty\\ \psi(Q)\geq cQ^{-\tau}\end{subarray}}K\cap A_{Q}(cQ^{-\tau},\mathbf{0})\in\mathscr{G}_{\mathrm{mea}}^{s}(K).$
Since the set on the left is contained in $K\cap W_{d}(\psi,\mathbf{0})$, we have
$K\cap W_{d}(\psi,\mathbf{0})\in\mathscr{G}_{\mathrm{mea}}^{s}(K),$
and hence
$\mathcal{H}^{s}\big(K\cap W_{d}(\psi,\mathbf{0})\big)=\mathcal{H}^{s}(K)=\infty,$
which completes the proof of (1.6) in Theorem 1.2. ∎
5. Proof of Corollary 1.3
(1) By (1.5) in Theorem 1.2, for any $1/d<\tau<\alpha$, we have
$\operatorname{\dim_{H}}\big(K\cap W_{d}(\psi_{\tau},{\bm{\theta}})\big)\\ \geq\min\biggl\{\delta-\frac{(\tau d-1)(\delta+\frac{1+d}{1+\tau}-d)}{\beta\delta},\ \delta+\frac{1+d}{1+\tau}-d\biggr\}.$
As $\tau\downarrow 1/d$, the right-hand side converges to $\delta$. Consequently,
$\lim_{\tau\downarrow 1/d}\operatorname{\dim_{H}}\big(K\cap W_{d}(\psi_{\tau},{\bm{\theta}})\big)\geq\delta.$
Since the reverse inequality is immediate, it then follows that
$\operatorname{\dim_{H}}\big(K\cap\operatorname{VWA}_{d}({\bm{\theta}})\big)=\delta.$
(2) Note that the assumptions $d=1$, $\theta=0$ and $\beta>0$ imply that (1.6) in Theorem 1.2 can be applied to obtain the Hausdorff measure of $K\cap W_{1}(\tau,0)$. Observe that
$E_{1}(\tau,0)\supset W_{1}(\psi_{\tau},0)\Big\backslash\bigcup_{n=1}^{\infty}W_{1}(\psi_{\tau+1/n},0).$
By the upper bound (1.4) in Theorem 1.2, for each $n\in\mathbb{N}$,
(5.1) $\operatorname{\dim_{H}}\big(K\cap W_{1}(\psi_{\tau+1/n},0)\big)\leq\delta+\frac{2}{1+\tau+1/n}-1<\delta+\frac{2}{1+\tau}-1.$
On the other hand, by (1.6) in Theorem 1.2,
$\mathcal{H}^{s}\big(K\cap W_{1}(\psi_{\tau},0)\big)=\infty,$
where
$s=\min\bigg\{\delta-\frac{(\tau-1)(1-\delta)(\delta+\frac{2}{1+\tau}-1)}{\beta\delta},\ \delta+\frac{2}{1+\tau}-1\bigg\}.$
A straightforward calculation shows that the minimum is $\delta+\frac{2}{1+\tau}-1$ whenever
$\tau>\frac{1-\delta^{2}-\beta\delta}{(1-\delta)^{2}}.$
By our assumption $\delta\geq 1-\beta/2$, we have $\beta\geq 2-2\delta$, and hence
$\frac{1-\delta^{2}-\beta\delta}{(1-\delta)^{2}}\leq\frac{1-\delta^{2}-(2-2\delta)\delta}{(1-\delta)^{2}}=1,$
which implies that $s=\delta+\frac{2}{1+\tau}-1$. It follows that
[…]
6. Some Examples
6.1. Irreducible self-similar sets and proof of Theorem 1.4
Fix a finite set $\Lambda$. An iterated function system (IFS for short) is a finite collection
$\mathcal{F}=\{\phi_{i}:i\in\Lambda\}$
of contractive similarities on $\mathbb{R}^{d}$; that is, for each $i\in\Lambda$, the map $\phi_{i}$ has the form
$\phi_{i}(\mathbf{x})=\rho_{i}O_{i}\mathbf{x}+\mathbf{b}_{i},$
where $0<\rho_{i}<1$, $O_{i}\in\operatorname{SO}_{d}(\mathbb{R})$ is an orthogonal matrix, and $\mathbf{b}_{i}\in\mathbb{R}^{d}$. It is shown in [17] that there exists a unique non-empty compact set $K=K(\mathcal{F})\subset\mathbb{R}^{d}$, called the self-similar set associated with $\mathcal{F}$, such that
(6.1) $K=\bigcup_{i\in\Lambda}\phi_{i}(K).$
Following [17], we say that $\mathcal{F}$ satisfies the open set condition (OSC for short) if there exists a non-empty bounded open set $U\subset\mathbb{R}^{d}$ such that
$\bigcup_{i\in\Lambda}\phi_{i}(U)\subset U\quad\text{and}\quad\phi_{i}(U)\cap\phi_{j}(U)=\emptyset\quad\text{for all }i\neq j.$
If $\mathcal{F}$ is an IFS satisfying the OSC, then the Hausdorff dimension of $K$ is given by $\delta$, which is the unique solution to
$\sum_{i\in\Lambda}\rho_{i}^{\delta}=1,$
and moreover,
$0<\mathcal{H}^{\delta}(K)<\infty.$
Let $\mu$ denote the normalized restriction of $\mathcal{H}^{\delta}$ to $K$, which we refer to as the self-similar measure. Then $\mu$ is $\delta$-Ahlfors regular and invariant under $\mathcal{F}$ in the sense that
(6.2) $\mu=\sum_{i\in\Lambda}\rho_{i}^{\delta}\cdot(\mu\circ\phi_{i}^{-1}).$
Let $\Lambda^{*}=\bigcup_{k\geq 1}\Lambda^{k}$ denote the set of finite words over $\Lambda$. Let $k\geq 1$. Given a word ${\bm{\omega}}=(\omega_{1},\dots,\omega_{k})\in\Lambda^{k}$, we define
(6.3) $\begin{split}&\phi_{\bm{\omega}}:=\phi_{\omega_{1}}\circ\cdots\circ\phi_{\omega_{k}},\quad\rho_{\bm{\omega}}:=\rho_{{\bm{\omega}}_{1}}\cdots\rho_{{\bm{\omega}}_{k}},\\ &K_{\bm{\omega}}:=\phi_{\bm{\omega}}(K)\quad\text{and}\quad\mu_{\bm{\omega}}:=\frac{\mu|_{K_{\bm{\omega}}}}{\mu(K_{\bm{\omega}})}=\mu\circ\phi_{\bm{\omega}}^{-1}.\end{split}$
We call $\mu_{\bm{\omega}}$ a branch of $\mu$. Intuitively, $\mu_{\bm{\omega}}$ represents the measure $\mu$ restricted to a ‘cylinder’ set $K_{\bm{\omega}}$ and then rescaled to a probability measure.
Finally, an IFS $\mathcal{F}$ is called strongly irreducible if $\mathbb{R}^{d}$ is the only finite union of affine subspaces of $\mathbb{R}^{d}$ that is invariant under all $\phi_{i}\in\mathcal{F}$. This condition ensures that the system cannot be confined to a lower-dimensional structure and is crucial for certain rigidity and dimension results.
The proof of the local estimate (see Corollary 6.5 below) for rational numbers near strong irreducible self-similar set relies on the equidistribution of self-similar measure in the homogeneous space $X=\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$, which has been proved in [3, 2, 20].
Set $\mathfrak{g}:=\mathfrak{sl}_{d+1}(\mathbb{R})$ to be the Lie algebra of $G=\operatorname{SL}_{d+1}(\mathbb{R})$. We equip $G$ with the unique right $G$-invariant Riemannian metric which coincides with $\|\cdot\|_{\mathfrak{g}}$ at $\mathfrak{g}=T_{\mathrm{Id}}G$. Write
$\mathcal{A}=\{E_{i,j}:1\leq i,j\leq d+1,\ i\neq j\}\cup\{E_{i,i}-E_{i+1,i+1}:i=1,\dots,d\}$
for the standard basis of $\mathfrak{g}$. Given $l\in\mathbb{N}$, write $\Xi_{l}$ for the set of words of length $l$ with letters in $\mathcal{A}$. Each $D\in\Xi_{l}$ acts as a differential operator on the space of smooth functions $C^{\infty}(X)$. Let $m_{X}$ be the $\operatorname{SL}_{d+1}(\mathbb{R})$-invariant Haar measure on $X$. Given $f\in C^{\infty}(X)$ and $p\in[1,\infty]$, we set
$\mathcal{S}_{p,l}(f)=\sum_{D\in\Xi_{l}}\|Df\|_{L^{p}},$
where $\|\cdot\|_{L^{p}}$ refers to the $L^{p}$-norm for $m_{X}$ on $X$. Let
$B^{\infty}_{p,l}(X)=\{f\in C^{\infty}(X):\mathcal{S}_{p,l}(f)<\infty\}.$
For $t>0,\mathbf{x}\in\mathbb{R}^{d}$, let
(6.4) $g(t)=\begin{pmatrix}e^{t/d}\mathbb{I}_{d}&0\\ 0&e^{-t}\end{pmatrix}\text{ and }u(\mathbf{x})=\begin{pmatrix}\mathbb{I}_{d}&\mathbf{x}\\ 0&1\end{pmatrix}$,
where $\mathbb{I}_{d}$ is the $d\times d$ identity matrix.
Let $\exp(\cdot)$ denote the exponential map from $\mathfrak{g}$ to $G$. Denote by $B_{r}^{\mathfrak{g}}(0)$ the ball in $\mathfrak{g}$ centered at 0 with radius $r$, and let $r_{0}$ be such that $\exp:B^{\mathfrak{g}}_{r_{0}}(0)\rightarrow\mathrm{exp}(B_{r_{0}}^{\mathfrak{g}}(0))$ is a diffeomorphism. For $y\in X$, we define the injectivity radius at $y$ as
$\mathrm{inj}(y)=\sup\{r\leq r_{0}:\text{ the map }g\mapsto gy\text{ is injective on }\exp(B_{r}^{\mathfrak{g}}(0))\}.$
This definition of injectivity radius is different from that in [3, 2], but this will not affect the final result. Next, we recall the following key equidistribution result from [2], which generalizes earlier special cases considered in [3, 20]. We remark that their result applies to a more general class of self-similar measures than the one defined in (6.2). However, such a level of generality is not needed for our purposes, and we therefore do not present it in full generality here.
Theorem 6.1 ([2, Theorem 1.2]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. There exists a constant $\kappa>0$ such that for all $t>0$, $y\in X$, and $f\in B^{\infty}_{p,l}(X)$, the following holds
$\bigg|\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu(\mathbf{x})-\int_{X}f{\,\mathrm{d}}m_{X}\bigg|\ll\mathrm{inj}(y)^{-1}\mathcal{S}_{p,l}(f)e^{-\kappa t},$
where the implicit constant only depends on the IFS $\mathcal{F}$ and $l=\lceil\frac{d(d+1)}{4}\rceil$. Moreover, the exponent $\kappa$ can be chosen so that it is nondecreasing as $\operatorname{\dim_{H}}K$ increases.
Remark 9.
The “Moreover” part is not explicitly stated in their paper, but it can be verified to hold. In fact, $\kappa$ is very small, while the proof of the zero–one law (See Theorem BHZ) only relies on its existence. This is one reason why the authors of [3, 2] did not explicitly derive $\kappa$.
As a corollary of the above theorem, we obtain an effective equidistribution result for the branches of $\mu$. The proof is based on the approach of Chen [10] in the one-dimensional setting and is included for completeness.
Corollary 6.2 ([10, Corollary 2.6]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Let $f\in B_{p,l}^{\infty}(X)$. Then, for any ${\bm{\omega}}\in\Lambda^{*}$ and $y\in X$, we have
$\bigg|\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\bm{\omega}}(\mathbf{x})-\int_{X}f{\,\mathrm{d}}m_{X}\bigg|\ll\mathrm{inj}(y)^{-1}\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}\mathcal{S}_{p,l}(f)e^{-\kappa t},$
where $\kappa$ is as in Theorem 6.1 and the implicit constant only depends on the IFS $\mathcal{F}$.
Proof.
Since $\mu_{\bm{\omega}}=\mu\circ\phi_{\bm{\omega}}^{-1}$ and $\phi_{\bm{\omega}}(\mathbf{x})=\rho_{\bm{\omega}}\mathbf{x}+\mathbf{b}_{\bm{\omega}}$ with $\mathbf{b}_{\bm{\omega}}=\phi_{\bm{\omega}}(\bf 0)$, we have
$\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\bm{\omega}}(\mathbf{x})$
[…]
$=$ $\int_{\mathbb{R}^{d}}f\left(g\left(t+\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{x})g\left(-\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{b}_{\bm{\omega}})y\right){\,\mathrm{d}}\mu(\mathbf{x}),$
where the last equality follows from
[…]
Therefore, by Theorem 6.1,
$\bigg|\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\bm{\omega}}(\mathbf{x})-\int_{X}f{\,\mathrm{d}}m_{X}\bigg|$
[…]
By a standard estimate, there exists a constant $C_{1}>0$ such that for any $h\in G$ and $z\in X$,
$\mathrm{inj}(hz)\geq C_{1}\,\|\mathrm{Ad}(h^{-1})\|_{\infty}\,\mathrm{inj}(z),$
where $\mathrm{Ad}$ denotes the adjoint action of $G$ on $\mathfrak{g}$ and $\|\cdot\|_{\infty}$ is the operator norm. Applying this with $h=g(-\frac{d}{d+1}\ln\rho_{\bm{\omega}})u(\mathbf{b}_{\bm{\omega}})$ yields
$\begin{split}\mathrm{inj}\left(g\left(-\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{b}_{\bm{\omega}})y\right)&\geq C_{1}\left\|\operatorname{Ad}\left(g\left(\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)\right)\right\|_{\infty}\mathrm{inj}(u(\mathbf{b}_{\bm{\omega}})y)\\ &\geq C_{1}\cdot\rho_{\bm{\omega}}^{-1}\,\mathrm{inj}(u(\mathbf{b}_{\bm{\omega}})y).\end{split}$
Since $u(\mathbf{b}_{\bm{\omega}})$ is uniformly bounded over ${\bm{\omega}}\in\Lambda^{*}$, there exists a constant $C_{2}>0$ such that $\mathrm{inj}(u(\mathbf{b}_{\bm{\omega}})y)\geq C_{2}\,\mathrm{inj}(y)$. Combining these bounds, we obtain
$\mathrm{inj}\left(g\left(-\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{b}_{\bm{\omega}})y\right)^{-1}\leq C_{1}^{-1}C_{2}^{-1}\rho_{\bm{\omega}}^{-1}\mathrm{inj}(y)^{-1},$
which immediately implies the desired effective equidistribution result. ∎
By approximating $A_{Q}(\eta,\mathbf{0})$ with smooth functions and applying the above effective equidistribution result, Khalil and Luethi [20] obtained the following estimate, which is stated in a slightly different form from their original result for our purposes.
Lemma 6.3 ([20, Theorem 9.1 and Lemma 12.7]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Let $Q\geq 1$ be a large integer and $0<\eta\leq Q^{-1/d}$. Then, there exists a constant $C_{\mathcal{F}}$ depending only on $\mathcal{F}$ such that
$Q\eta^{d}-C_{\mathcal{F}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}}\ll\mu\left(A_{Q}(\eta,\bf 0)\right)\ll Q\eta^{d}+C_{\mathcal{F}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}},$
where $\kappa$ is as in Theorem 6.1 and the implied constants are independent of $\mathcal{F}$.
Remark 10.
In fact, Khalil and Luethi [20] proved a result stronger than that stated in Lemma 6.3. Let
$\begin{split}E_{Q}(\eta,\mathbf{0}):=\bigg\{\mathbf{x}\in\mathbb{R}^{d}:\bigg|\mathbf{x}-\frac{\mathbf{p}}{q}\bigg|<\frac{\eta}{q}&\text{ for some $\frac{\mathbf{p}}{q}\in\mathbb{Q}^{d}$}\\ &\text{with $0<q<2Q$ and $\gcd(q,\mathbf{p})=1$}\bigg\}.\end{split}$
Clearly,
$E_{Q}(\eta,\mathbf{0})\setminus E_{Q/2}(\eta,\mathbf{0})\subset A_{Q}(\eta,\mathbf{0})\subset E_{Q}(\eta,\mathbf{0}).$
The upper and lower bounds given in Lemma 6.3 are obtained precisely by estimating the $\mu$-measures of these two outer sets, as respectively given in [20, Theorem 9.1 and Lemma 12.7].
Motivated by this result together with Corollary 6.2, Chen [10] derived the following estimate for $\mu_{\bm{\omega}}(A_{Q}(\eta,\bf 0))$. Although Chen’s formulation is somewhat different from ours, it essentially implies the same result.
Proposition 6.4 ([10, Proposition 2.8]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Let $Q\geq 1$ be a large integer and $0<\eta\leq Q^{-1/d}$. Then, there exists a constant $C_{\mathcal{F}}$ depending only on $\mathcal{F}$ such that for any ${\bm{\omega}}\in\Lambda^{*}$,
$Q\eta^{d}-C_{\mathcal{F}}\,\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}}\ll\mu_{\bm{\omega}}\big(A_{Q}(\eta,{\bf 0})\big)\ll Q\eta^{d}+C_{\mathcal{F}}\,\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}},$
where $\kappa$ is as in Theorem 6.1 and the implied constants are independent of $\mathcal{F}$.
The main term $Q\eta^{d}$ in Proposition 6.4 dominates provided that
$Q\eta^{d}\geq 2C_{\mathcal{F}}\,\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}(Q^{-1}\eta)^{\frac{d\kappa}{d+1}}\quad\Longleftrightarrow\quad\rho_{\bm{\omega}}\geq 2C_{\mathcal{F}}\,Q^{-1}\eta^{-(d-\frac{d\kappa}{d+1})/(1+\frac{d\kappa}{d+1})}.$
In particular, if $\eta\geq Q^{-\alpha}$ with $\alpha>1/d$, then the above inequalities hold if
(6.5) $\rho_{\bm{\omega}}\geq 2C_{\mathcal{F}}\,Q^{-1}Q^{\alpha(d-\frac{d\kappa}{d+1})/(1+\frac{d\kappa}{d+1})}=2C_{\mathcal{F}}\,Q^{-(1+\frac{d\alpha(d+1)\kappa}{d+1+d\kappa}-d\alpha)}.$
Therefore, Proposition 6.4 yields the desired local estimate for rational points in a neighbourhood of self-similar sets.
Corollary 6.5.
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Choose $\alpha$ and $\beta$ so that
$\alpha>1/d\mathrm{\quad and\quad}0<\beta<1+\frac{d\alpha(d+1)\kappa}{d+1+d\kappa}-d\alpha,$
where $\kappa$ is as in Theorem 6.1. Then, $\mu$ satisfies the $(\alpha,\beta,\mathbf{0})$-local counting property.
Proof.
Since $K$ satisfies the open set condition, any ball $B$ contains a branch $K_{{\bm{\omega}}}$ of $K$ with $|B|\asymp|K_{{\bm{\omega}}}|$, and is itself contained in a finite union of branches of $K$, namely
$B\subset\bigcup_{i=1}^{i_{B}}K_{{\bm{\omega}}_{i}},$
where $i_{B}\asymp 1$ and $|B|\asymp|K_{{\bm{\omega}}_{i}}|$ for all $i$. By the self-simlarity and the Ahlfors regularity of $\mu$, we have
$|B|\asymp|K_{\bm{\omega}}|\quad\Longrightarrow\quad\mu(B)\asymp\mu(K_{\bm{\omega}}).$
Therefore, the corollary follows if, for any ${\bm{\omega}}\in\Lambda^{*}$ with $\rho_{\bm{\omega}}\geq Q^{-\beta}$ and $Q^{-\alpha}\leq\eta\leq Q^{-1/d}$,
$\mu\big(K_{\bm{\omega}}\cap A_{Q}(\eta,{\bf 0})\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta^{d}.$
This is indeed the case, since by Proposition 6.4 and (6.5),
[…]
Proof of Theorem 1.4.
(1) By Corollary 6.5, the $\delta$-Ahlfors regular self-similar measure $\mu$ supported on $K$ satisfies the $(\alpha,\beta,\mathbf{0})$-local counting property, where
$\alpha>1/d\mathrm{\quad and\quad}0<\beta<1+\frac{d\alpha(d+1)\kappa}{d+1+d\kappa}-d\alpha,$
and $\kappa$ is as in Theorem 6.1. The conclusion is an immediate consequence of Corollaries 1.3 (1).
(2) Suppose that $d=1$ and $\theta=0$. Denote by $\kappa(K)$ the exponent in Theorem 6.1 associated to the self-similar set $K$. For any $\kappa_{0}>0$, choose $\alpha=\alpha(\kappa_{0})>1$ such that
$\beta:=1+\frac{2\alpha\kappa_{0}}{2+\kappa_{0}}-\alpha>0.$
This is always possible, since as $\alpha\downarrow 1$ the left-hand side converges to $\frac{2\kappa_{0}}{2+\kappa_{0}}>0$. Suppose that $\kappa_{0}$ is small enough so that there are nontrivial self-similar sets $K$ satisfying $\kappa(K)>\kappa_{0}$. Let $0<\delta_{0}<1$ be such that $\delta_{0}>1-\beta_{0}/2$ and $\kappa(K)>\kappa_{0}$ whenever $\operatorname{\dim_{H}}K\geq\delta_{0}$. The second restriction is possible because, by Theorem 6.1, the parameter $\kappa(K)$ is non-decreasing as the Hausdorff dimension of $K$ increases.
Let $K\subset\mathbb{R}$ be a self-similar set with $\operatorname{\dim_{H}}K=\delta\geq\delta_{0}$. By the definition of $\delta_{0}$, we have $\kappa(K)>\kappa_{0}$, and so
$1+\frac{2\alpha\kappa(K)}{2+\kappa(K)}-\alpha>1+\frac{2\alpha\kappa_{0}}{2+\kappa_{0}}-\alpha=\beta.$
By Corollary 6.5, $\mu$ satisfies the $(\alpha,\beta,0)$-local counting property with $\alpha$ and $\beta$ as defined above. Since $\delta\geq\delta_{0}\geq 1-\beta/2$, it follows from Corollary 1.3 (2) that for any $1<\tau<\alpha$,
$\mathcal{H}^{\delta+\frac{2}{1+\tau}-1}\big(K\cap E_{1}(\tau,0)\big)=\infty.\qed$
6.2. Missing digits sets and proof of Theorem 1.5
In [31], Yu employed Fourier-analysis methods to show that sufficiently thick missing digits measures supported on $\mathbb{R}$ satisfy the $(\alpha,\beta,0)$-local counting property. We begin by recalling some relevant definitions.
Let $b>2$ be an integer, and let
$\mathcal{D}\subset\{0,1,\dots,b-1\}^{d}$
be a non-empty proper subset. The associated missing digits set is defined by
$K(b,\mathcal{D}):=\left\{\mathbf{x}\in[0,1]^{d}:\mathbf{x}=\sum_{n=1}^{\infty}\frac{\mathbf{d}_{n}}{b^{n}},\ \mathbf{d}_{n}\in\mathcal{D}\text{ for all }n\in\mathbb{N}\right\}.$
Let $\mu_{b,\mathcal{D}}$ be the natrual normalized Hausdorff measure supported on $K(b,\mathcal{D})$, which will be referred to as the missing digits measure.
For a probability measure $\mu$ supported on $\mathbb{R}^{d}$, its Fourier $l^{1}$-dimension is defined by
$\dim_{l^{1}}\mu:=\sup\left\{s\geq 0:\sum_{\bm{\xi}\in\mathbb{Z}^{d}:|\bm{\xi}|\leq R}|\widehat{\mu}(\bm{\xi})|\ll R^{d-s}\text{ for all $R>0$}\right\},$
where
$\widehat{\mu}(\bm{\xi}):=\int_{\mathbb{R}^{d}}e^{-2\pi i\bm{\xi}\cdot\mathbf{x}}{\,\mathrm{d}}\mu(\mathbf{x})$
is the Fourier transformation of $\mu$ and $\bm{\xi}\cdot\mathbf{x}$ denotes the inner product.
Let
$A(q,\eta,{\bm{\theta}}):=\{\mathbf{x}\in\mathbb{R}^{d}:\|q\mathbf{x}-{\bm{\theta}}\|<\eta\}.$
In [31, §6.2], Yu proved that for some $\varepsilon>0$ (which can be chosen arbitrarily small but, once fixed, remains constant), as long as $Q$ is sufficiently large and
$1>\eta\geq Q^{1-1/(d-\dim_{l^{1}}\mu)+\varepsilon},$
then
(6.6) $\sum_{q=Q}^{2Q}\mu\big(A(q,\eta,{\bm{\theta}})\big)\asymp Q\eta^{d}.$
Since we will require that $\eta<Q^{-1/d}$, the above estimate is meaningful only when
$1-\frac{1}{d-\dim_{l^{1}}\mu}<-1/d\quad\Longrightarrow\quad\dim_{l^{1}}\mu>d-\frac{d}{d+1}.$
The above analysis indicates that for measures $\mu$ with sufficiently large $\dim_{l^{1}}\mu$, one has a good counting estimate for rational points in neighborhoods of the corresponding fractal set. However, determining $\dim_{l^{1}}\mu$ is in general a highly nontrivial problem. In [31, 32], Yu calculated $\dim_{l^{1}}\mu_{b,\mathcal{D}}$ for the missing digits measure and in particular showed that for any $t\geq 1$,
$\liminf_{\begin{subarray}{c}b\to\infty\\ \#\mathcal{D}\geq b^{d}-t\end{subarray}}\dim_{l^{1}}\mu_{b,\mathcal{D}}=d.$
This establishes the existence of nontrivial fractal measures with arbitrarily large Fourier $l^{1}$-dimension.
From now on, we restrict our attention to the missing digits sets in $\mathbb{R}$, and for simplicity we denote $K(b,\mathcal{D})$ and $\mu_{b,\mathcal{D}}$ by $K$ and $\mu$, respectively. The restriction to one dimension is simply because, at present, we are unable to establish the local counting property for missing digits measures in $\mathbb{R}^{d}$ with $d\geq 2$.
For any ${\bm{\omega}}\in\mathcal{D}^{*}:=\bigcup_{k\geq 1}\mathcal{D}^{k}$, denote by $K_{\bm{\omega}}$ and $\mu_{\bm{\omega}}$ the branch of $K$ and $\mu$ associated to $\bm{\omega}$}, repectively (see (6.3) for precise definitions). To prove a local version of (6.6), observe that for any ${\bm{\omega}}\in\mathcal{D}^{*}$,
(6.7) $|\widehat{\mu_{\bm{\omega}}}(\xi)|=|\widehat{\mu}(b^{-|{\bm{\omega}}|}\xi)|,$
where $|{\bm{\omega}}|$ denotes the length of $\bm{\omega}$}. Consequently, it is shown in [31, Pages 40-41] that for any $1/2<s_{1}<\dim_{l^{1}}\mu$, and any $\alpha$ and $\beta$ satisfying $1<\alpha<\frac{s_{1}}{1-s_{1}}$ and $0<\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}$, we have
(6.8) $\sum_{q=Q}^{2Q}\mu_{\bm{\omega}}\big(A(q,\eta,\theta)\big)\asymp Q\eta$
whenever $Q$ is large enough (depending on $\alpha$ and $\beta$ only), $\theta\in\mathbb{R}$, $Q^{-\alpha}\leq\eta\leq Q^{-1}$ and $|K_{\bm{\omega}}|\geq Q^{-\beta}$. Note that $\mu_{\bm{\omega}}=\frac{\mu|_{K_{\bm{\omega}}}}{\mu(K_{\bm{\omega}})}$ for all ${\bm{\omega}}\in\mathcal{D}^{*}$ (see (6.3)). It follows that
(6.9) $\sum_{q=Q}^{2Q}\mu\big(K_{\bm{\omega}}\cap A(q,\eta,\theta)\big)=\mu(K_{\bm{\omega}})\cdot\sum_{q=Q}^{2Q}\mu_{\bm{\omega}}\big(A(q,\eta,\theta)\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta.$
Although $A_{Q}(\eta,\theta)=\bigcup_{q=Q}^{2Q}A(q,\eta,\theta)$, this does not imply that
$\mu\big(K_{\bm{\omega}}\cap A_{Q}(\eta,\theta)\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta$
because the sets $A(q,\eta,\theta)$ may have overlap. In [31, Pages 57-58], after a carefully analysis of the measure of the overlaps, Yu was able to establish the above result for $\theta=0$. For arbitrary $\theta\in\mathbb{R}$, although some of his ideas are applicable, there are still several nontrivial difficulties.
Lemma 6.6.
Suppose that $\dim_{l^{1}}\mu>1/2$. For any $1/2<s_{1}<\dim_{l^{1}}\mu$, let $1<\alpha<\frac{s_{1}}{1-s_{1}}$ and $0<\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}$. Then, for any $\theta\in\mathbb{R}$, $\mu$ satisfies the $(\alpha,\beta,\theta)$-local counting property.
Proof.
To prove the lemma, it is enough to show that
(6.10) $\mu\big(K_{\bm{\omega}}\cap A_{Q}(\eta,\theta)\big)\asymp\mu(K_{\bm{\omega}})\cdot Q\eta,$
whenever $Q$ is sufficiently large (depending only on $\alpha$ and $\beta$), $\theta\in\mathbb{R}$, $Q^{-\alpha}\leq\eta\leq Q^{-1}$, and $|K_{\bm{\omega}}|\geq Q^{-\beta}$.
The upper bound of (6.10) follows directly from (6.9), so it remains to establish the corresponding lower bound. Define the function
$f_{Q}=\sum_{q=Q}^{2Q}\mathbbm{1}_{A(q,\eta,\theta)}.$
Note that $f_{Q}$ is a non-negative integer-valued function. If $f_{Q}(x)=k>0$, then $x\in A(q,\eta,\theta)$ for precisely $k$ integers $q$ with $Q\leq q\leq 2Q$. In particular, $x\in A_{Q}(\eta,\theta)$. Therefore,
$0<f_{Q}(x)<k\quad$ $\Longrightarrow\quad f_{Q}(x)/k<1=\mathbbm{1}_{A_{Q}(\eta,\theta)}(x)$
(6.11) $\Longrightarrow\quad\frac{\int_{\{f_{Q}<k\}}f_{Q}{\,\mathrm{d}}\mu_{\bm{\omega}}}{k}\leq\mu_{\bm{\omega}}\big(A_{Q}(\eta,\theta)\big).$
We claim that there exists $\rho\in(0,1)$ such that for any $2\leq k<Q$,
(6.12) $\int_{\{f_{Q}\geq k\}}f_{Q}{\,\mathrm{d}}\mu_{\bm{\omega}}\ll\frac{Q\eta}{k}+Q^{\rho}\eta\log_{2}Q,$
where the implied constant is independent of $Q,k,\theta$ and $\eta$. This gives that for $k$ and $Q$ large enough,
[…]
where the last inequality follows from
[…]
whenever $k$ and $Q$ are large enough.
Now, we prove the claim, and in the course of the proof it will become clear how the parameters $\alpha,\beta,s_{1}$ in the lemma arise. Let $Q^{-\alpha}\leq\eta<Q^{-1}/4$. Once the lower bound in (6.10) is established for this case, the corresponding lower bound for $\eta$ in the range $Q^{-1}/4\leq\eta\leq Q^{-1}$ follows immediately. Let $x\in\mathbb{R}$. If $f_{Q}(x)=m\geq 2$, then there exist exactly distinct $Q\leq q_{1}<\cdots<q_{m}\leq 2Q$ such that
$\|q_{i}x-\theta\|<\eta\quad\text{for $1\leq i\leq m$}.$
Consequently, for $1\leq i\leq m-1$
$\|(q_{m}-q_{i})x\|=\|q_{m}x-\theta-(q_{i}x-\theta)\|\leq 2\eta.$
This implies that $\|qx\|<2\eta$ has m1m-1 solutions for $q\leq Q$. Let $q_{\max}$ be the largest integer for which $\|q_{\max}x\|<2\eta$, and let $a/n$ be the reduced fraction of $p_{\max}/q_{\max}$. Since the distance between any two distinct rationals $p/q$ and $p^{\prime}/q^{\prime}$ is at least $1/(qq^{\prime})$, we see that if $\|qx\|<2\eta,\|q^{\prime}x\|<2\eta$ with $q>q^{\prime}$, then
(6.13) $\bigg|x-\frac{p}{q}\bigg|<\frac{2\eta}{q}\text{ and }\bigg|x-\frac{p^{\prime}}{q^{\prime}}\bigg|<\frac{2\eta}{q^{\prime}}\quad\Longrightarrow\quad\bigg|\frac{p}{q}-\frac{p^{\prime}}{q^{\prime}}\bigg|<\frac{4\eta}{q^{\prime}}.$
Since $\eta<Q^{-1}/4$, the last inequality is less than $1/(qq^{\prime})$, and so $p/q=p^{\prime}/q^{\prime}$. Therefore, any rational number $p/q$ satisfying $|x-p/q|<2\eta/q$ has the form $\frac{at}{nt}$ for some integer $t\geq 1$. Since there are m1m-1 many such rationals $p/q$ and all the denumerator is less than $Q$, we have that $1\leq n\leq q_{\max}/(m-1)\leq Q/(m-1)$. Hence,
(6.14) $\bigg|x-\frac{a}{n}\bigg|=\bigg|x-\frac{p_{\max}}{q_{\max}}\bigg|<\frac{2\eta}{q_{\max}}<\frac{2\eta}{n(m-1)}\quad\Longrightarrow\quad\|nx\|<\frac{2\eta}{m-1}.$
In summary, we have proved
$f_{Q}(x)=m\quad\Longrightarrow\quad\|nx\|<\frac{2\eta}{m-1}\text{ for some $n\leq\frac{Q}{m-1}$}.$
This implies that
$\int_{\{f_{Q}\in[m,2m)\}}f_{Q}{\,\mathrm{d}}\mu_{\bm{\omega}}$ $\leq 2m\int\mathbbm{1}_{\bigcup_{q=1}^{Q/(m-1)}A(q,2\eta/(m-1),0)}{\,\mathrm{d}}\mu_{\bm{\omega}}$
(6.15) $\leq 2m\int\sum_{q=1}^{Q/(m-1)}\mathbbm{1}_{A(q,2\eta/(m-1),0)}{\,\mathrm{d}}\mu_{\bm{\omega}}.$
Define a new function
(6.16) $g=\sum_{q=1}^{Q/(m-1)}\mathbbm{1}_{A(q,2\eta/(m-1),0)}.$
We can modify $g$ by replacing the charater functions with smooth functions just as in the proof of [31, Theorem 4.1], and we see that it is possible to show that
(6.17) $\int_{[0,1]}g{\,\mathrm{d}}\mu_{\bm{\omega}}\ll\int_{[0,1]}g(x){\,\mathrm{d}}x+\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\bm{\omega}}}(\xi)\widehat{g}(\xi)|.$
The first term on the right can be easily computed and we have
[…]
For the second term, observe that for $\xi\neq 0$, we have
$|\widehat{g}(\xi)|\ll\sum_{q=1}^{Q/(m-1)}\eta/m\cdot\mathbbm{1}_{q|\xi}\leq d(\xi)\eta/m,$
where $d(\xi)$ denote the number of divisors of $\xi$. The proof of this inequality can be found in [31, page 16]. This together with (6.7) gives
$\begin{split}\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\bm{\omega}}}(\xi)\widehat{g}(\xi)|&\ll\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu_{\bm{\omega}}}(\xi)|d(\xi)\eta/m\\ &=\eta/m\cdot\sum_{\xi\in\mathbb{Z}:0<|\xi|\leq Q/\eta}|\widehat{\mu}(b^{-|{\bm{\omega}}|}\xi)|d(\xi).\end{split}$
It then follows from an argument in [31, Page 40] that for each $1/2<s_{1}<\dim_{l^{1}}\mu,$
$\sum_{0<|\xi|\leq Q/\eta}|\widehat{\mu}(b^{-|{\bm{\omega}}|}\xi)|d(\xi)\ll b^{s_{1}|{\bm{\omega}}|}Q^{1-s_{1}}\eta^{-(1-s_{1})}.$
Therefore, for
(6.19) $1<\alpha<\frac{s_{1}}{1-s_{1}}\mathrm{\quad and\quad}0<\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}},$
we have that
(6.20) $b^{s_{1}|{\bm{\omega}}|}Q^{1-s_{1}}\eta^{-(1-s_{1})}\leq Q^{\rho}\quad\text{with $\rho=s_{1}\beta+(1+\alpha)(1-s_{1})\in(0,1)$}$
whenever $Q$ is large enough (depending on $\alpha$ and $\beta$ only), $Q^{-\alpha}\leq\eta\leq Q^{-1}$ and $|K_{\bm{\omega}}|=b^{-|{\bm{\omega}}|}\geq Q^{-\beta}$. This means that
[…]
Note that
$\int_{\{f_{Q}\geq k\}}f_{Q}\,{\,\mathrm{d}}\mu_{\bm{\omega}}=\sum_{j\geq 0:\,2^{j}k\leq Q}\int_{\{f_{Q}\in[2^{j}k,2^{j+1}k)\}}f_{Q}\,{\,\mathrm{d}}\mu_{\bm{\omega}}.$
Since there are at most $\log_{2}Q$ possibilities for $j$, it follows from (6.21) that
$\begin{split}\sum_{j\geq 0:\,2^{j}k\leq Q^{1-\rho^{\prime}}}\int_{\{f_{Q}\in[2^{j}k,2^{j+1}k)\}}f_{Q}\,{\,\mathrm{d}}\mu_{\bm{\omega}}&\ll\sum_{j\geq 0:\,2^{j}k\leq Q}\biggl(\frac{Q\eta}{2^{j}k}+Q^{\rho}\eta\biggr)\\ &\ll\frac{Q\eta}{k}+Q^{\rho}\eta\log_{2}Q,\end{split}$
which completes the proof of the claim. ∎
Remark 11.
The choice of the parameters $\alpha$ and $\beta$ in Lemma 6.6 indeed satisfies
$\alpha\leq\frac{1+\delta-\beta\delta}{1-\delta},$
as stated in Remark 1. To prove this inequality, it suffices to show that
$1-\alpha+(1+\alpha-\beta)\delta\geq 0.$
Using the upper bound $\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}$, we obtain
$\begin{split}1-\alpha+(1+\alpha-\beta)\delta&\geq 1-\alpha+\bigg(1+\alpha-\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}\bigg)\delta\\ &=1+\alpha(\delta/s_{1}-1).\end{split}$
The desired conclusion then follows from the fact that, for missing digits measures, $\delta\geq\dim_{l^{1}}\mu>s_{1}$ (see [31, Lemma 1.4 (1) and §3.6]).
Proof of Theorem 1.5.
(1) Since $\dim_{l^{1}}\mu>1/2$, it follows from Lemma 6.6 that $\mu$ satisfies the $(\alpha,\beta,\mathbf{0})$-local counting property for every $\theta\in\mathbb{R}$. The conclusion follows directly from Corollary 1.3 (1).
(2) Since $\operatorname{\dim_{H}}K\dim_{l^{1}}\mu=\delta\cdot\dim_{l^{1}}\mu>1/2$, we can choose $s_{1}$ such that $1/2<s_{1}<\dim_{l^{1}}\mu$ and $\delta s_{1}>1/2$. With this choice of $s_{1}$, we further choose parameters $\alpha$ and $\beta$ such that
(6.22) $1<\alpha<\frac{2\delta s_{1}-s_{1}}{1-s_{1}}\mathrm{\quad and\quad}2(1-\delta)\leq\beta<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}}.$
In order for the set of admissible $\beta$ values to be nonempty, and thus for $\delta\geq 1-\beta/2$ to hold, the upper bound that $\alpha$ can take must be less than $\frac{s_{1}}{1-s_{1}}$. Indeed, the set of admissible $\beta$ values is nonempty if and only if
$2(1-\delta)<\frac{1-(1+\alpha)(1-s_{1})}{s_{1}},$
which is exactly equivalent to
$\alpha<\frac{2\delta s_{1}-s_{1}}{1-s_{1}},$
as stated in (6.22). Once $\beta$ is chosen according to the lower bound in (6.22), the inequality $\delta\geq 1-\beta/2$ is automatically satisfied.
For the parameters $\alpha$ and $\beta$ chosen above, it follows from Lemma 6.6 that $\mu$ satisfies the $(\alpha,\beta,0)$-local counting property. Since $\delta\geq 1-\beta/2$, by Corollary 1.3 (2), we have
$\mathcal{H}^{\delta+\frac{2}{1+\tau}-1}\big(K\cap E_{1}(\tau,0)\big)=\infty,$
whenever $1<\tau<\alpha$. ∎
Acknowledgements
Y. He was supported by the NSFC (No. 12401108) and partially by a grant from the Guangdong Provincial Department of Education (2025KCXTD013).